- Status Report
- August 14, 2022
NOAA GOES SATOPS Morning Report: Friday, February 18, 2005
OPS 49 rev 26790 07 at BOSS-D 14:05z Lost. Unable to connect via SOCC DCC or ESOC DCC. EPH loads and Data recovery delayed until next rev. Cat 2 declared: 37 PBK sec/148 S/C seconds of TF data lost due to Site change of next rev. Next rev has an earlier acquisition that doesn’t allow PR2 to complete it’s record. All smooth data recovered on R+9.
All other DMSP operations were nominal over the past 24 hours.
All POES operations were nominal over the past 24 hours.
GOES spacecraft have started battery reconditioning in preparation for the upcoming eclipse periods.
All GOES operations were nominal over the past 24 hours.
Questions or comments to Keith.Amburgey@noaa.gov